OpenStudy (anonymous):

Evaluate and tell if it is extraneous or not. x^2-6x+13=0

OpenStudy (anonymous):

@geny55

OpenStudy (anonymous):

Step 1 : Raise x to the 2nd power

OpenStudy (anonymous):

factor?

OpenStudy (anonymous):

this is the Equation at the end of step 1 : (x2 - 6x) + 13 = 0

OpenStudy (anonymous):

Step 2 : Simplify x2-6x + 13

OpenStudy (anonymous):

3+ or - 2i

OpenStudy (anonymous):

Equation at the end of step 2 : x2 - 6x + 13 = 0

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

so what would the answer be? bc i factored it

OpenStudy (anonymous):

Simplifying x2 + 6x + 13 = 0

OpenStudy (anonymous):

Reorder the terms: 13 + 6x + x2 = 0 Solving 13 + 6x + x2 = 0 Solving for variable 'x'. Begin completing the square. Move the constant term to the right: Add '-13' to each side of the equation. 13 + 6x + -13 + x2 = 0 + -13 Reorder the terms: 13 + -13 + 6x + x2 = 0 + -13 Combine like terms: 13 + -13 = 0 0 + 6x + x2 = 0 + -13 6x + x2 = 0 + -13 Combine like terms: 0 + -13 = -13 6x + x2 = -13 The x term is 6x. Take half its coefficient (3). Square it (9) and add it to both sides. Add '9' to each side of the equation. 6x + 9 + x2 = -13 + 9 Reorder the terms: 9 + 6x + x2 = -13 + 9 Combine like terms: -13 + 9 = -4 9 + 6x + x2 = -4 Factor a perfect square on the left side: (x + 3)(x + 3) = -4 Can't calculate square root of the right side. The solution to this equation could not be determined.